coffee price $2.00 morning quantity 240 afternoon quantity 70 price $1.50 morning quantity 266 afternoon quantity...

coffee

price $2.00 morning quantity 240 afternoon quantity 70

price $1.50 morning quantity 266 afternoon quantity 123

price elasticity?

NPV?
Answer #1

We can see a decrease in price.

Change in price = 2-1.50 = 0.50

% change in price = (0.50 / 2) * 100 = 25%

Quantity demanded increased by 26 in the morning.

% change in quantity demanded = (26 / 240 ) * 100 = 10.83%

Quantity demanded increased by 53 in the afternoon.

% change in quantity demanded = (53 / 70) * 100 = 75.71%

price elasticity (in the morning) = (% change in quantity demanded ) / (% change in price )

= 10.83% / 25% = 0.43

The value is less than 1 indicating it is inelastic.

price elasticity (in the afternoon) = (% change in quantity demanded ) / (% change in price )

= 75.71% / 25%

= 3.02

price elasticity of demand is greater than 1 which means it is elastic. That means more customers are willing to drink coffee.